OpenStudy (liziekay):

Write the following in terms of sin θ and cos θ; then simplify if possible. (Leave your answer in terms of sin θ and/or cos θ.) sin θ cot θ + 8 cos θ

OpenStudy (liziekay):

@zepdrix

zepdrix (zepdrix):

Eyyyy Lisa May. Remember how to turn cotangent into sines and cosines?

OpenStudy (liziekay):

Well, I know that tan = sin theta/ cos theta

zepdrix (zepdrix):

k good. cotangent is the flip of tangent. so cot = cos / sin

zepdrix (zepdrix):

\[\large\rm \sin \theta\color{orangered}{\cot \theta}+8\cos \theta\]becomes,\[\large\rm \sin \theta\color{orangered}{\frac{\cos \theta}{\sin \theta}}+8\cos \theta\]That's a good first step, ya?

OpenStudy (liziekay):

agreed

zepdrix (zepdrix):

Hmm what next? Any ideas?

OpenStudy (misssmartiez):

That awkward moment when a person asks something that you have no clue about. *cringes* I'll go with zepdrix. I don't really exceed much in math. I'd stick to L-arts, English, and science for now.

OpenStudy (liziekay):

I don't think we can elliminate any variables

OpenStudy (liziekay):

but sin is 1/csc theta so maybe that helps?

zepdrix (zepdrix):

No? Remember than when you have something sitting on it's own, it's the same as having it in the numerator,\[\large\rm a\frac{b}{c}=\frac{ab}{c}\]

zepdrix (zepdrix):

No, we've sines and cosines, let's leave it that way. We have a cancellation somewhere though. Not seeing it yet?

OpenStudy (liziekay):

right! I totally blanked out. You could elliminate sin theta!

OpenStudy (liziekay):

would it be 9 cos since you have two of them?

zepdrix (zepdrix):

1cos +8cos left over after cancellation? And combining like-terms? Yayy good job!

OpenStudy (liziekay):

That was quite simple, I was way over thinking it!!